Fermi National Laboratory
 

Operation Summary for May 23, 2011 - May 25, 2011

 

Monday, May 23

The day shift began with Operations monitoring the Tevatron (TEV) store 8760 (91.58E30) and the Recycler stash (136.9E10), with the Antiproton Source (Pbar) stacking (79.4E10), with the Main Injector (MI), Linac, and Booster operating normally, with the Recycler Pelletron having communication problems, with NuMI off due to target replacement, with the MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Pelletron personnel continued investigating the Electron Cooling (ECool) communication problems. TEV personnel began crystal collimator studies at 10:15 AM. FESS personnel reported that they had repaired the TEV sector B1 LCW pump electrical breaker and returned the pump to service. Linac personnel switched from the I- Source to the H- Source. Crystal collimator studies ended at 4 PM.

During the evening shift, Operations terminated store 8760 at 4:25 PM. TEV technicians performed necessary maintenance repairs. At 5:51 PM, Pelletron personnel reported that ECool communication could not be restored and that the lack of electron cooling would result in lower luminosity stores. Operations put the TEV into HEP shot setup at 6:20 PM and established store 8762 at 7:30 PM with an average initial luminosity (AIL) of 114.69E30.

Tuesday May 24

The midnight shift began with Operations monitoring store 8762 (59.58E30) and the stash (146.3E10), with Pbar stacking (120.6E10), and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8762 at 3:35 AM, put the TEV into shot setup, and established store 8763 at 4:45 AM with an AIL of 126.33E30.

During the day shift, Pelletron personnel continued their ECool communications investigation. TEV personnel conducted an end-of-store study (EOS) from 10 AM to noon. Linac personnel switched from the H- Source to the I- Source. Operations terminated store 8762 at 12:15 PM. Operations loaded an antiproton-only store for further TEV studies.

During the evening shift, Pelletron personnel reported that they had restored ECool communications at 5 PM. Operators searched and secured MI-31. The TEV EOS studies concluded at 7:50 PM and Operations put the TEV into shot setup. Operations established store 8765 at 11:50 PM with an AIL of 350.79E30.

Wednesday May 25

The midnight shift began with Operations monitoring store 8765 (291.78E30) and the stash (95.9E10), with Pbar stacking (18.1E10), and with MiniBooNE taking beam.

During the midnight shift, operators reported that a MI RF station (MIRF4) tripped off on a tunnel cavity LCW leak indication; operators turned off the station.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow TEV and Booster studies.
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/25/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory